OpenStudy (anonymous):

Find a Taylor series for cos x centered at x=pi/4.

OpenStudy (anonymous):

The Taylor series (centered at \(x=c\)) for cosine is given by \[\cos x=\sum_{n=0}^\infty\frac{(-1)^n(x-c)^{2n}}{(2n)!}\]